OpenStudy (anonymous):

Help.......The function f(x) = 15(2)x represents the growth of a frog population every year in a remote swamp.

OpenStudy (anonymous):

The function f(x) = 15(2)x represents the growth of a frog population every year in a remote swamp. Elizabeth wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Elizabeth's purposes?

OpenStudy (anonymous):

f(x) = 15(2 to the one half power)2x f(x) = 15(2)the x over 2 power f(x) = fifteen halves(2)x f(x) = 30(2)x

ganeshie8 (ganeshie8):

\(\large 15(2)^x = 15(2^{\frac{1}{2}})^{2x}\)

OpenStudy (anonymous):

Thanks!:)

ganeshie8 (ganeshie8):

yw !